solve for y
y+3x=8

Answer by jim_thompson5910(35256)

Subtract 3x from both sides to get

y + 3x = 8

y + 3x - 3x = 8 - 3x

y + 0x = 8 - 3x

y = 8 - 3x

y = -3x + 8


Final Answer: y = -3x + 8